# DiffOG: Differentiable Policy Trajectory Optimization with Generalizability

### v1

@@ Abstract
Imitation learning-based visuomotor policies excel at manipulation tasks but often produce suboptimal action trajectories compared to model-based methods. Directly mapping camera data to actions via neural networks can result in jerky motions and difficulties in meeting critical constraints, compromising safety and robustness in real-world deployment. For tasks that require high robustness or strict adherence to constraints, ensuring trajectory quality is crucial. However, the lack of interpretability in neural networks makes it challenging to generate constraint-compliant actions in a controlled manner. This paper introduces differentiable policy trajectory optimization with generalizability (DiffOG), a learning-based trajectory optimization framework designed to enhance visuomotor policies. By leveraging the proposed differentiable formulation of trajectory optimization with transformer, DiffOG seamlessly integrates policies with a generalizable optimization layer. Visuomotor policies enhanced by DiffOG generate smoother, constraint-compliant action trajectories in a more interpretable way. DiffOG exhibits strong generalization capabilities and high flexibility. We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies while having minimal impact on policy performance, outperforming trajectory processing baselines such as greedy constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy.
@@ III-A Action Space Assumption
The proposed method assumes that the discrete-time derivative $\frac{d{{c}}_{t}}{dt}$ in (3) represents decoupled explicit accelerations or velocities.
Specifically, among commonly used action spaces for robot policies, Assumption 1 holds under the following conditions:
@@
2. The action space consists of the robot end-effector velocity, defined as the linear velocity in $xyz$ combined with the angular velocity in Euler angles $[{\phi},{\theta},{\psi}]^{\mathrm{T}}$, which can be denoted as $[v_{\text{xyz}},\dot{\phi},\dot{\theta},\dot{\psi}]^{\mathrm{T}}\in\mathbb{R}^{6}$. $v_{\text{xyz}}$ represents the liner velocity of the end-effector. In this case, the derivative of the robot end-effector velocity corresponds to the linear acceleration combined with how the Euler angular velocity components change over time, which can be demoted as $[\dot{v}_{\text{xyz}},\ddot{\phi},\ddot{\theta},\ddot{\psi}]^{\mathrm{T}}\in\mathbb{R}^{6}$. This action space is commonly used in imitation learning for manipulation, such as [12, 26, 36, 32].
@@ III-C Differentiable Trajectory Optimization
Proposition 1 guarantees that optimization problem (6) is feasible and strictly convex, ensuring the existence of a unique optimal solution $\mathbf{y}^{\star}_{t}$. However, optimization problem (6) constitutes a meaningful trajectory optimization only under Assumption 1. This is because, under Assumption 1, the constraints in (7) correspond to physically meaningful velocity or acceleration bounds, and the smoothing term $(\frac{d\hat{\mathbf{c}}_{t}}{dt})^{\mathrm{T}}\frac{d\hat{\mathbf{c}}_{t}}{dt}$ represents a physically interpretable quantity which reflects the trajectory’s smoothness. For instance, a large $(\frac{d\hat{\mathbf{c}}_{t}}{dt})^{\mathrm{T}}\frac{d\hat{\mathbf{c}}_{t}}{dt}$ indicates a lack of smoothness in the trajectory. Consequently, optimization problem (6) can enforce trajectory smoothing by minimizing the associated cost. Moreover, Assumption 1 ensures that the condition $d_{\text{min}}<d_{\text{max}}$ carries explicit physical meaning. For example, if the action space is defined in terms of joint angles, this condition ensures that the minimum and maximum joint velocity constraints do not conflict. Only when both Assumption 1 and the condition $d_{\text{min}}<d_{\text{max}}$ hold does the constraining in optimization problem (6) remain valid and physically meaningful.
@@ V Experiments
Figure 3: We validated DiffOG on 13 different tasks. The 13 tasks span three types, corresponding to the three action spaces described in Assumption 1.
@@ V-A Benchmark with Baselines Across Tasks
In this section, we demonstrate how DiffOG optimizes the trajectories generated by base policies across 13 tasks. The 13 tasks span three types, corresponding to the three action spaces described in Assumption 1. For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [12]. For the end-effector linear position action space, we include six tasks: push-T from [5], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [33]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [34]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [35].
@@
1. In the simulation benchmarks, we evaluate trajectory smoothness by analyzing the first-order derivatives of actions. Specifically, we compare the maximum values and standard deviations of these derivatives throughout the policy rollouts. In Robomimic, the first-order derivative corresponds to angular acceleration in Euler angles, while in Meta-World and push-T, it represents linear velocity. For the real-world tasks, in addition to computing the maximum and standard deviation of the first-order derivatives (i.e., joint angular velocities), we also assess the second-order derivatives (joint angular accelerations) using the same metrics.
@@ VII-A Hyperparameters and Dataset
The bound of 0.1 for Robomimic tasks represents normalized acceleration. To convert the it back to the raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for rotational acceleration of Euler angles [12]. For example, the maximum linear acceleration is $0.1\times 20~{}\text{m}/\text{s}^{2}=2~{}\text{m}/\text{s}^{2}$.
@@ VII-B Computation of Trajectory Metrics
The acceleration metrics for Robomimic tasks are normalized values, and are averaged on linear accelerations and rotational acceleration of Euler angles. Please refer to Section VII-A to see more information of unnormalizing them to their raw form, which are measured in $\text{m}/\text{s}^{2}$ and $\text{rad}/\text{s}^{2}$.

### v3

@@ Abstract
Imitation learning-based visuomotor policies excel at manipulation tasks but often produce suboptimal action trajectories compared to model-based methods. Directly mapping camera data to actions via neural networks can result in jerky motions and difficulties in meeting critical constraints, compromising safety and robustness in real-world deployment. For tasks that require high robustness or strict adherence to constraints, ensuring trajectory quality is crucial. However, the lack of interpretability in neural networks makes it challenging to generate constraint-compliant actions in a controlled manner. This paper introduces differentiable policy trajectory optimization with generalizability (DiffOG), a learning-based trajectory optimization framework designed to enhance visuomotor policies. By leveraging the proposed differentiable formulation of trajectory optimization with transformer, DiffOG seamlessly integrates policies with a generalizable optimization layer. DiffOG refines action trajectories to be smoother and more constraint-compliant while maintaining alignment with the original demonstration distribution, thus avoiding degradation in policy performance. We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies while having minimal impact on policy performance, outperforming trajectory processing baselines such as greedy constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy. Please visit the project website for more details: https://zhengtongxu.github.io/diffog-website/.
@@ III-A Action Space Assumption
The proposed method assumes that the discrete-time derivative $\frac{d{{c}}_{t}}{dt}$ in (3) represents explicit, element‑wise accelerations or velocities.
The purpose of introducing Assumption 1 is to justify the addition of constraints in subsequent sections. The fundamental question addressed here is whether imposing constraints on $\frac{d{{c}}_{t}}{dt}$ can physically restrict the actions generated by the policy. When Assumption 1 holds, the answer is straightforward: the constraints directly regulate explicit physical quantities such as velocities or accelerations. However, even when Assumption 1 does not strictly hold, certain forms of constraints may still provide meaningful regulation over the robot’s motion. We elaborate on these cases in the following discussion.
Among commonly used action spaces for robot policies, Assumption 1 holds under the following conditions:
@@
Another widely adopted policy action space is $[v_{x},\,v_{y},\,v_{z},\,\dot{r}_{x},\,\dot{r}_{y},\,\dot{r}_{z}]^{\mathrm{T}}\in\mathbb{R}^{6},$ where $\mathbf{v}_{xyz}=[v_{x},v_{y},v_{z}]^{\mathrm{T}}$ denotes the translational velocity, and $\dot{\mathbf{r}}=[\dot{r}_{x},\dot{r}_{y},\dot{r}_{z}]^{\mathrm{T}}$ is the time derivative of the axis–angle rotation vector [12, 26, 36, 32]. Taking the time derivative gives $[\dot{v}_{x},\,\dot{v}_{y},\,\dot{v}_{z},\,\ddot{r}_{x},\,\ddot{r}_{y},\,\ddot{r}_{z}]^{\!\mathrm{T}}.$ The first three elements are explicit linear accelerations, whereas the last three are axis-angle second derivatives, not the physical angular acceleration $\dot{\bm{\omega}}$. Therefore, this type of action space is not covered by Assumption 1. However, we will show that under this type of action representation, constraints on the action derivative can still effectively regulate the robot’s physical motion.
The relationship between the axis-angle representation and the angular acceleration is governed by a Jacobian matrix $\mathbf{J}(\mathbf{r})$, which maps $\dot{\mathbf{r}}$ to the angular velocity $\bm{\omega}$ such that $\bm{\omega}=\mathbf{J}(\mathbf{r})\dot{\mathbf{r}}$. Taking the time derivative, the angular acceleration is given by $\dot{\bm{\omega}}=\mathbf{J}(\mathbf{r})\ddot{\mathbf{r}}+\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$, where $\ddot{\mathbf{r}}$ is the second derivative of the axis-angle vector and $\dot{\mathbf{J}}(\mathbf{r})$ is the time derivative of the Jacobian.
Although $\ddot{\mathbf{r}}$ does not directly represent $\dot{\bm{\omega}}$, constraining $\ddot{\mathbf{r}}$ still effectively regulates $\dot{\bm{\omega}}$. In the ideal case of small rotation angles and slow rotational motions, the term $\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$ becomes negligible, and $\mathbf{J}(\mathbf{r})$ approximates the identity matrix. This yields $\dot{\bm{\omega}}\approx\ddot{\mathbf{r}}$, making constraints on $\ddot{\mathbf{r}}$ nearly equivalent to constraining the true angular acceleration.
In practice, while $\mathbf{J}(\mathbf{r})$ may deviate from the identity matrix, it typically remains bounded. This boundedness is evidenced by the widespread empirical success of using the axis-angle velocity as the rotational action in many imitation learning frameworks, such as those in [12, 26, 36]. These works employ end-to-end learning without any explicit constraints on $\dot{\mathbf{r}}$ and $\ddot{\mathbf{r}}$, yet still achieve rotationally stable behaviors. This empirical stability indicates that the mapping from axis-angle velocity to true angular velocity is well-behaved, further supporting the validity of this action space.
Building on this insight, DiffOG explicitly constrains and smooths $\ddot{\mathbf{r}}$ (Section III-C), enabling finer control over rotational motion and resulting in smoother, more physically consistent behaviors. Therefore, even though this action representation does not strictly satisfy Assumption 1, it still enables effective regularization of angular acceleration. Experimental results in Section V-A confirm that DiffOG led to improved motion for this type of action space.
@@ III-C Differentiable Trajectory Optimization
Proposition 1 guarantees that optimization problem (6) is feasible and strictly convex, ensuring the existence of a unique optimal solution $\mathbf{y}^{\star}_{t}$. Moreover, the smoothing term $(\frac{d\hat{\mathbf{c}}_{t}}{dt})^{\mathrm{T}}\frac{d\hat{\mathbf{c}}_{t}}{dt}$ represents a physically interpretable quantity which reflects the trajectory’s smoothness. For instance, a large $(\frac{d\hat{\mathbf{c}}_{t}}{dt})^{\mathrm{T}}\frac{d\hat{\mathbf{c}}_{t}}{dt}$ indicates a lack of smoothness in the trajectory. Consequently, optimization problem (6) can enforce trajectory smoothing by minimizing the associated cost. Finally, the condition $d_{\text{min}}<d_{\text{max}}$ carries explicit physical meaning. For example, if the action space is defined in terms of joint angles, this condition ensures that the minimum and maximum joint velocity constraints do not conflict.
@@ V Experiments
Figure 3: We validated DiffOG on 13 different tasks. The 13 tasks span three types, corresponding to the three action spaces discussed in Section III-A.
@@ V-A Benchmark with Baselines Across Tasks
In this section, we demonstrate how DiffOG optimizes the trajectories generated by base policies across 13 tasks. The 13 tasks span three types, corresponding to the three action spaces described in Section III-A. For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [12]. For the end-effector linear position action space, we include six tasks: push-T from [5], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [33]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [34]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [35].
@@
1. In the simulation benchmarks, we evaluate trajectory smoothness by analyzing the first-order derivatives of actions. Specifically, we compare the maximum values and standard deviations of these derivatives throughout the policy rollouts. In Robomimic, the first-order derivative corresponds to the second derivative of axis-angle, while in Meta-World and push-T, it represents linear velocity. For the real-world tasks, in addition to computing the maximum and standard deviation of the first-order derivatives (i.e., joint angular velocities), we also assess the second-order derivatives (joint angular accelerations) using the same metrics.
@@ VII-A Hyperparameters and Dataset
The bound of 0.1 for Robomimic tasks represents normalized acceleration. To convert the it back to the raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for second derivative of axis-angle [12]. For example, the maximum linear acceleration is $0.1\times 20~{}\text{m}/\text{s}^{2}=2~{}\text{m}/\text{s}^{2}$.
@@ VII-B Computation of Trajectory Metrics
The acceleration metrics for Robomimic tasks are normalized values, and are averaged on linear accelerations and second derivatives of axis-angle. Please refer to Section VII-A to see more information of unnormalizing them to their raw form, which are measured in $\text{m}/\text{s}^{2}$ and $\text{rad}/\text{s}^{2}$.